OpenStudy (anonymous):

Simplify the expression. (2/3) * (15.5 + (-24) ) + (-18)

OpenStudy (anonymous):

A. 48 B. 52 C. 16 D. 22

OpenStudy (anonymous):

that doent help

OpenStudy (anonymous):

*dont

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@NeonStrawsForever

OpenStudy (phi):

can you do this part 15.5 + (-24) ?

OpenStudy (phi):

I would do "normal subtraction" 24 - 15.5 and then make the answer negative

OpenStudy (anonymous):

-8.5

OpenStudy (phi):

ok so (2/3) * (15.5 + (-24) ) + (-18) becomes (2/3) * -8.5 -18 order of operations: do multiply first what is 2*-8.5 ?

OpenStudy (anonymous):

no its 2/3 the fraction

OpenStudy (phi):

yes. but 2/3 means (or at least you can interpret it as) multiply by 2, then divide by 3 btw, after we are done , none of the choices match the number we get. Are you sure there is no typo?

OpenStudy (anonymous):

no thoes are th choices to choose from

OpenStudy (anonymous):

@phi its -17

OpenStudy (phi):

(2 * -8.5)/3 = -17/3 now we have (2/3) * (15.5 + (-24) ) + (-18) (2/3) * -8.5 - 18 -17/3 - 18 we can change 18 to 3*18/3 so the problem is \[ - \frac{17}{3} - \frac{54}{3} \]

OpenStudy (anonymous):

im not good with fractions i suck at them

OpenStudy (phi):

that becomes \[ - \frac{71}{3} = -23 \frac{2}{3} \]

OpenStudy (anonymous):

ok

OpenStudy (phi):

can you post a screen shot of the original question? Something is wrong.